### v1

# Matrix product state approach to lossy boson sampling and noisy IQP sampling

Abstract
Sampling problems have emerged as a central avenue for demonstrating quantum advantage on noisy intermediate-scale quantum devices. However, physical noise can fundamentally alter their computational complexity, often making them classically tractable. Motivated by the recent success of matrix product state (MPS)-based classical simulation of Gaussian boson sampling (Oh et al., 2024), we extend this framework to investigate the classical simulability of other noisy quantum sampling models. We develop MPS-based classical algorithms for lossy boson sampling and noisy instantaneous quantum polynomial-time (IQP) sampling, both of which retain the tunable accuracy characteristic of the MPS approach through the bond dimension. Our approach constructs pure-state decompositions of noisy or lossy input states whose components remain weakly entangled after circuit evolution, thereby providing a means to systematically explore the boundary between quantum-hard and classically-simulable regimes. For boson sampling, we analyze single-photon, Fock, and cat-state inputs, showing that classical simulability emerges at transmission rates scaling as $O(1/\sqrt{N})$, reaching the known boundary of quantum advantage with a tunable and scalable method. Beyond reproducing previous thresholds, our algorithm offers significantly improved control over the accuracy–efficiency trade-off. It further extends the applicability of MPS-based simulation to broader classes of noisy quantum sampling models, including IQP circuits.
preprint: APS/123-QED
I Introduction
It is widely believed that a quantum computer can efficiently solve certain problems that are hard for classical computers, such as integer factoring and Hamiltonian simulation [1, 2]. Since one of the most critical obstacles to realizing a quantum computer is noise, quantum error correction (QEC) has been developed to actively detect and correct errors, enabling fault-tolerant quantum computation [3, 4]. Despite substantial advances in both quantum hardware and the theory and implementation of QEC [5, 6], currently available quantum devices cannot yet fully exploit the QEC technique in a fully scalable manner. Consequently, extensive efforts have focused on achieving quantum advantage using currently available quantum devices, often referred to as noisy intermediate-scale quantum (NISQ) devices [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17].
To pursue quantum advantage with NISQ devices, quantum sampling problems have emerged as promising candidates due to their complexity-theoretic hardness results in noiseless cases and their relatively feasible experimental requirements, compared with more sophisticated quantum algorithms [7, 8, 9]. The most representative examples of sampling problems include boson sampling, random circuit sampling, and instantaneous quantum polynomial-time (IQP) sampling [7, 18, 19]. In fact, the experimental implementation of these sampling problems has been conducted to claim quantum advantage by experiments [20, 12, 14, 9, 21, 17].
However, such claims have been continuously challenged due to the noise in the realistic experiment (e.g., Refs. [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]). The main insight for this is that physical noise often degrades and ultimately destroys quantum advantage. For example, when a depolarizing noise occurs in quantum circuits and the circuit depth is large, the system accumulates entropy, causing the output state to converge to the maximally mixed state, which is easily simulated classically [36]. Similarly, in photonic systems, it has been shown that photon loss makes a quantum state converge to a thermal state, which is easy to classically simulate [25, 26]. As such, physical noise may make sampling problems that are hard to classically simulate when noiseless become easy to simulate. Hence, to rigorously assess the potential advantage of quantum devices, it is essential to understand whether and when sampling problems become easy to classically simulate under realistic noise conditions.
Focusing specifically on boson sampling, one of the most widely studied sampling models, as aforementioned, when the loss rate is sufficiently large, boson sampling becomes classically easy. Motivated by this observation, a variety of classical simulation algorithms have been developed [37, 25, 38, 39, 27, 32]. One line of work [25] used a strategy to approximate a lossy quantum state by a quantum state that is classically easy to simulate. For example, Ref. [25] and Ref. [24] showed that a lossy single photon state can be approximated by a thermal state or a separable state, respectively, when the loss rate is high. Furthermore, this idea has been extended to lossy Gaussian boson sampling (GBS) [26]. Nonetheless, when applied to recent GBS experiments, while it provides an important benchmark, the classical algorithm was not sufficient to simulate them [14, 12]. A key limitation of these approaches is that, since they rely on approximating a given state by the nearest classically simulable state, their performance cannot be systematically improved by allocating additional computational resources. Hence, below a loss rate threshold, the algorithm fails to operate.
Notably, a more recently proposed classical algorithm based on the matrix product state (MPS) has overcome this limitation and successfully simulated the state-of-the-art GBS experiments at the moment. The crucial feature of the new classical algorithm is the tunability of its performance by providing more computational resources, which translates to the bond dimension of the MPS. However, because the method in Ref. [33] relies on the Gaussianity of the quantum system, it has not been extended to other systems so far, despite its high performance in practice.
In this work, we generalize the MPS for GBS proposed in Ref. [33] to simulate lossy boson sampling and noisy IQP sampling. In other words, we develop MPS-based classical algorithms for lossy boson sampling and noisy IQP sampling that correspond to the state-of-the-art classical algorithm for lossy GBS; hence, our classical algorithms also enjoy the property that the accuracy is tunable by consuming computational resources, unlike the previous algorithms [25, 24]. The central idea of our approach is to construct a pure-state decomposition of a lossy or noisy input state such that each component remains weakly entangled after circuit evolution. For the boson sampling case, we consider various input states, such as single-photon states, Fock states, and cat states as input and analyze the performance of our algorithm using the entanglement entropy [40, 41]. Notably, we find a similar classically simulable range of the transmission rate to GBS and previous results [33, 24, 25], which is given as $O(1/\sqrt{N})$, where $N$ is the input photon number. We then estimate the bond dimension required to simulate, providing guidelines for future experiments. We further extend our framework to noisy IQP sampling and then demonstrate that the performance of our classical algorithm is comparable to that of the recently proposed algorithm [42].
The remainder of this paper is organized as follows. In Sec. II, we introduce the problem setup by describing boson sampling and IQP sampling circuits and provide the relation between MPS and efficient classical simulation in Sec. III. In Sec. IV, we present an MPS-based classical simulation algorithm for boson sampling with various input states and analyze the asymptotically simulable range and obtain the numerical estimation of computational resources. In Sec. V, we provide a classical simulation algorithm for IQP sampling together with its numerical analysis. Finally, Sec. VI discusses the implications of our results and concludes the paper.
II Problem setup
II.1 Boson sampling
Boson sampling is a sampling problem that is proposed to demonstrate quantum advantage using photons [7]. To formally introduce boson sampling, consider an $M$-mode bosonic system and let $\hat{a}_{i}$ and $\hat{a}_{i}^{\dagger}$ denote the bosonic annihilation and creation operators of the $i$th mode, respectively, satisfying the canonical commutation relations, $\hat{a}_{i},\hat{a}_{j}^{\dagger}]=\delta_{ij}$ and $\hat{a}_{i},\hat{a}_{j}]=0$, where $1\leq i\leq M$. The input state of boson sampling is $N$ single photons in the first $N$ modes out of a total of $M$ modes, with the rest of the modes being initialized as vacuum, i.e., the input state is written as
$|1\rangle^{\otimes N}|0\rangle^{\otimes(M-N)}=\left(\prod_{i=1}^{N}\hat{a}_{i}^{\dagger}\right)|0\rangle^{\otimes M}.$ (1)
The boson sampling circuit is given by a linear optical circuit, characterized by an $M\times M$ unitary matrix $U$. More specifically, the latter transforms the input bosonic creation operators as
$\hat{a}_{j}^{\dagger}\rightarrow\hat{b}_{j}^{\dagger}\equiv\sum_{k=1}^{M}U_{jk}^{*}\hat{a}_{k}^{\dagger},$ (2)
where $\hat{b}_{i}^{\dagger}$ represents the creation operator of the $i$th output mode. Here, $U$ is often chosen to follow the Haar-random unitary ensemble for hardness results [7]. Such a Haar-random linear optical circuit can be implemented using two-mode beam splitters and phase shifters, with the circuit depth being linear in $M$ [43]. After the circuit, the output state is measured by photon number-resolving detectors, which give an output sample $(t_{1},\dots,t_{M})\in\mathbb{Z}_{\geq 0}^{M}$, where $t_{i}$ represents the output photon number on the $i$th output mode, satisfying $\sum_{i=1}^{M}t_{i}=N$ since linear-optical circuits preserve the photon number.
Boson sampling has attracted considerable attention due to its hardness result. More specifically, lossless boson sampling is proven to be hard to classically simulate under some plausible conjectures [7]. The main observation for the hardness result is that the corresponding output probability is described by a matrix function, so-called the permanent:
$p(t_{1},t_{2},\dots,t_{M})$ $=\dfrac{\left|\text{Perm }U_{T}\right|^{2}}{t_{1}!\cdots t_{M}!},$ (3)
which is known as #P-hard to compute approximately in multiplicative error. Here, $U_{T}$ is a matrix obtained by keeping the first $N$ columns from $U$ and repeating the $j$th row $t_{j}$ times.
Although the ideal boson sampling is believed to be hard to simulate classically, the actual experiments that implement boson sampling suffer from practical imperfections that hinder the demonstration of quantum advantage. In particular, one of the dominant imperfections is photon loss in the current experiments [44], and many studies demonstrate that under photon loss, photon sampling becomes easier to classically simulate [25, 24, 27, 26, 32, 30]. Therefore, it is crucial to consider the effect of photon loss and understand its significance for rigorously evaluating the classical simulability of experimental boson sampling.
In a lossy optical circuit, photon loss can be modeled using a beam splitter. More specifically, a loss channel on the $i$th mode with its loss rate $1-\eta$ (i.e., transmission rate $\eta$) transforms the $i$th mode’s bosonic creation operator as
$\hat{a}_{i}^{\dagger}\to\sqrt{\eta}\hat{a}_{i}^{\dagger}+\sqrt{1-\eta}\hat{e}_{i}^{\dagger},$ (4)
where $\hat{e}_{i}^{\dagger}$ is the creation operator of the corresponding environmental mode. One may easily check that a single-photon state transforms as follows under the loss channel of transmission rate $\eta$:
$|1\rangle\langle 1|\rightarrow(1-\eta)|0\rangle\langle 0|+\eta|1\rangle\langle 1|.$ (5)
Throughout this work, we assume that the optical devices have the same transmission rate and each mode has the same depth for simplicity; therefore, the transmission rate and the loss rate of each mode are assumed to be uniform. This uniform loss assumption allows each layer of the loss channel to commute with all other linear optical devices and be brought to the forefront of the lossless circuit. Consequently, the model reduces to a Haar random interferometer preceded by a uniform loss channel, as shown in Fig. 1. In the case of nonuniform loss, we can generalize this method by commuting the uniform loss part to the front and setting the overall transmission rate as the maximum among the modes in the remaining optics, and commuting to the front (see Ref. [24] for more details).
Figure 1: Photon loss in boson sampling circuit. (a) Lossy boson sampling circuit. Blue rectangles represent ideal beam splitters and phase shifters, while orange squares represent photon loss channels. (b) Assuming a uniform transmission rate, we separate the loss channel from the ideal linear optics and locate it in front of the ideal optics. $\eta$ is defined as the transmission rate of the total loss channel.
II.2 IQP sampling
IQP sampling is another sampling problem based on qubits, which is also widely believed to be computationally intractable for classical computers [19]. The $n$-qubit IQP sampling operates as follows. First, the input state is prepared as $|0\rangle^{\otimes n}$ and then evolves under a quantum circuit. At first, a layer of $n$ Hadamard gates is applied to the initial state $|0\rangle^{\otimes n}$, which transforms the state to the uniform superposition of all bitstrings with length $n$:
$\frac{1}{\sqrt{2^{n}}}\sum_{\vec{x}\in\{0,1\}^{n}}|\vec{x}\rangle.$ (6)
The qubits then go through the circuit with depth $d$ consisting of diagonal gates in $Z$ basis, such as $Z$, $S$, $T$, and $CZ$ gates. The diagonal gates transform the quantum state to
$\dfrac{1}{\sqrt{2^{n}}}\sum_{\vec{x}\in\{0,1\}^{n}}e^{if(\vec{x})}|\vec{x}\rangle,$ (7)
where $f$ is a real function determined by the diagonal gates. Thus, the diagonal gates change the relative phases between different computational basis elements. Finally, a layer of Hadamard gates is applied after diagonal gates, and the output state is measured in the $Z$ basis. Like boson sampling, IQP sampling is believed to be classically intractable: if it admitted an efficient classical simulation, the polynomial hierarchy would collapse to the third level, a consequence widely considered implausible [19].
As for the boson sampling case, however, many recent results show that IQP sampling may also be efficiently simulated when noise occurs in the system [45, 42, 46]. In this work, we exploit our MPS method to investigate the effect of noise on the hardness of IQP circuits and compare with other existing methods. To investigate the effect of noise on the complexity of simulating IQP sampling, we consider Pauli noises in IQP circuits, which is a dominant noise model in many physical devices. A Pauli noise channel transforms a quantum state $\hat{\rho}$ as follows:
[…]
$=(1-p_{X}-p_{Y}-p_{Z})\hat{\rho}+p_{X}\hat{X}\hat{\rho}\hat{X}+p_{Y}\hat{Y}\hat{\rho}\hat{Y}+p_{Z}\hat{Z}\hat{\rho}\hat{Z}$, (8)
where $p_{X}$, $p_{Y}$, and $p_{Z}$ are probabilities of $X$, $Y$, and $Z$ errors that characterize the noise rates of a Pauli channel. Our method applies to Pauli noise channels that can be decomposed into single-Pauli noise channels (i.e., channels involving only $X$, $Y$, or $Z$ errors). In what follows, we focus on two important special cases: dephasing noise and depolarizing noise. We first consider dephasing noise to provide an intuition, and then generalize it to the depolarizing noise. Here, the dephasing channel is a Pauli noise channel with $p_{X}=p_{Y}=0$, i.e., the noise effect is only by $Z$ error:
$\mathcal{N}_{0,0,p}(\hat{\rho})=(1-p)\hat{\rho}+p\hat{Z}\hat{\rho}\hat{Z}$, (9)
where dephasing noise rate $0\leq p\leq 1/2$. The dephasing noise channel is particularly simple to analyze in IQP sampling because it commutes with all diagonal gates in an IQP circuit, as both dephasing noise and the actions of diagonal gates are in the computational basis, which is similar to the loss channel in boson sampling. Hence, all dephasing noise channels can be moved forward, just after the first layer of Hadamard gates, by commuting with other gates. Assuming that the dephasing noise of the same noise rate is applied for each depth of gates for simplicity, the total noise channel applied to each input of the circuit when the noise channels are moved forward, as in Fig. 2(b), is given as a dephasing noise $\mathcal{N}_{0,0,p_{d}}$ with dephasing noise rate $p_{d}=(1-(1-2p)^{d})/2$.
Figure 2: Setup of a noisy IQP circuit. (a) The circuit consists of $Z$-diagonal gates (blue rectangles) with dephasing noise (orange squares) applied at each gate. Measurements are performed in the $X$ basis, equivalently implemented by a Hadamard layer followed by $Z$-basis readout. (b) Because dephasing commutes with $Z$-diagonal gates, all dephasing channels can be propagated to the input and consolidated as $\mathcal{N}_{0,0,p_{d}}$ acting before the circuit.
Generalizing this, we also investigate the classical simulation of a noisy IQP circuit under depolarizing noise:
$\mathcal{N}_{\text{depol}}(\hat{\rho})=\left(1-\dfrac{3}{2}p\right)\hat{\rho}+\dfrac{p}{2}\hat{X}\hat{\rho}\hat{X}+\dfrac{p}{2}\hat{Y}\hat{\rho}\hat{Y}+\dfrac{p}{2}\hat{Z}\hat{\rho}\hat{Z}$, (10)
where $0\leq p\leq 1/2$. After a qubit goes through this channel, the state is not changed when $p=0$ and becomes maximally mixed when $p=1/2$. Depolarizing noise can be expressed in terms of Pauli noise. Using the equality
[…]
we can derive $(1-q)^{3}+q^{3}=1-2p\Rightarrow q=(1-\sqrt{1-2p})/2$. For all range of $0\leq p\leq 1/2$, there exists corresponding $0\leq q\leq 1/2$.
Figure 3: Setup of a depolarizing-noisy IQP circuit. (a) The IQP circuit comprises $Z$-diagonal gates (blue rectangles), with a single-qubit depolarizing channel (orange squares) applied at each gate. Measurements are performed in the $X$ basis. (b) Each depolarizing channel is represented as a probabilistic mixture of Pauli errors $X$, $Y$, and $Z$, each occurring with rate $q$.
III MPS and classical simulability
As mentioned in the introduction, this work focuses on the MPS method for classical simulation. Thus, we provide a brief review of the MPS, which is particularly useful when a given quantum state is slightly entangled [41]. More specifically, consider a quantum state $|\psi\rangle$ written as
$|\psi\rangle=\sum_{i_{1},\dots,i_{L}=0}^{d-1}c_{i_{1},\cdots,i_{L}}|i_{1}\rangle\otimes\cdots\otimes|i_{L}\rangle,$ (12)
where each local Hilbert space dimension is $d$ and $\{|i_{k}\rangle\}_{i_{k}=0}^{d-1}$ represents an orthogonal basis of the $k$th local system. The MPS description of the state is a (approximate) decomposition of each amplitude as a product of matrices as
$c_{i_{1},\dots,i_{L}}\approx\sum_{\alpha_{1},\cdots,\alpha_{L-1}=0}^{\chi-1}\Gamma_{\alpha_{1}}^{[1]i_{1}}\lambda_{\alpha_{1}}^{[1]}\Gamma_{\alpha_{1}\alpha_{2}}^{[2]i_{2}}\lambda_{\alpha_{2}}^{[2]}\Gamma_{\alpha_{2}\alpha_{3}}^{[3]i_{3}}\cdots\Gamma_{\alpha_{L-1}}^{[L]i_{L}}.$ (13)
[…]
As the above indicates, the efficiency of the MPS description and the size of the required bond dimension are determined by the distribution of the Schmidt coefficients for each bipartition. More specifically, the scaling of the required bond dimension to describe a state is known to be characterized by the entanglement Rényi entropy (ERE) between the bipartitions of a pure state on $L$ sites $|\psi\rangle\in\mathcal{H}^{\otimes L}$. Here, the Rényi entropy is defined as
[…]
where $0\leq\alpha<\infty$ and $\alpha\neq 1$, and in the limit $\alpha\rightarrow 1$, it recovers the von Neumann entropy. The ERE of a pure state $|\psi\rangle$ between the subsystem $[1:l]$ and $[l+1:L]$ is defined as
$S_{\alpha}^{l}(|\psi\rangle)\equiv S_{\alpha}(\text{tr}_{[l+1:L]}(|\psi\rangle\langle\psi|)).$ (15)
In particular, $|\psi\rangle$ is classically approximable by MPS, i.e., the required bond dimension is at most polynomial in $L$, when $S_{\alpha}^{l}(|\psi\rangle)=O(\log L)$ for all $l$-site reduced states of $|\psi\rangle$ [41]. Using this as a sufficient simulability condition and relating it to the photon loss rate or Pauli noise rate, we can obtain the simulable range of loss or noise, respectively.
On the other hand, it is also well-known that if the ERE scales algebraically, i.e., $S_{\alpha}^{l}(|\psi\rangle)=\Omega(L^{\kappa})$ and $\kappa>0$ for some $l$, the corresponding state cannot be efficiently approximated by an MPS [41]. Using this, we also characterize the regime of loss or noise that is hard to simulate using MPS.
IV Classical simulation of lossy boson sampling
IV.1 MPS for simulating noisy quantum systems
As emphasized in Sec. III, the MPS is particularly useful for simulating low-entangled quantum systems. Therefore, the MPS method can be treated as a suitable choice to simulate noisy quantum systems because noise typically destroys the entanglement. However, an immediate obstacle to applying the MPS method for simulating noisy quantum systems, such as lossy boson sampling, is that the noisy quantum states are not pure in general, while the MPS method supports only pure states. Thus, the MPS method’s restriction to pure states hinders us from exploiting its feature for simulating noisy systems. An interesting way to overcome this limitation is to vectorize the density matrix so that we can describe a mixed state by a pure state formalism, which is called the MPO method [47, 27, 32]. Indeed, this approach exhibits that as the system becomes noisier, it is easier to perform classical simulation. However, compared to the MPS method, the MPO method requires more bond dimensions and more resources for updating the state, resulting in a greater resource requirement.
More recently, another method has been proposed to overcome this limitation, simulating state-of-the-art GBS experiments classically [33]. The main idea of this method is to decompose a noisy mixed state into a probabilistic mixture of pure states, each of which has low entanglement as the system becomes noisier. Such a method is the state-of-the-art classical simulator for GBS, to the best of our knowledge. Therefore, its generalization to other noisy quantum systems may promise to find one of the best classical algorithms at this moment. Nevertheless, the proposed decomposition appears to be highly dependent on the Gaussianity of the state, and thus, its generalization to more general cases has not been conducted. We now generalize this method to lossy boson sampling and noisy IQP sampling by finding an appropriate decomposition of noisy quantum states, each of which has low entanglement.
IV.2 Decomposition of lossy input state
Consider a lossy boson sampling circuit that consists of $N$ single-photon inputs in the first $N$ modes and MNM-N vacuum inputs, as introduced in Sec. II.1. After the input state in Fig. 1(b) passes through the loss channel, as in Fig. 4(a), it transforms to
$\hat{\rho}_{\text{in}}\equiv\hat{\sigma}^{\otimes N}\otimes|0\rangle\langle 0|^{\otimes(M-N)},$ (16)
where $\hat{\sigma}\equiv(1-\eta)|0\rangle\langle 0|+\eta|1\rangle\langle 1|$ is the lossy single-photon state. As highlighted before, since the resultant output state is no longer pure, it does not immediately allow an MPS description.
Instead of employing the MPO method, which requires substantial computational resources, we adopt the MPS formalism to simulate mixed states. In this approach, a mixed state is expressed as a convex combination of pure states, and each pure state is then simulated according to its corresponding probability weight [33]. More specifically, if we have a mixed state input $\hat{\rho}_{\text{in}}=\sum_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$ and the rest of the quantum circuit is unitary $\hat{U}$, one may simply sample $i$ from the probability distribution $\{p_{i}\}$ and then simulate $\hat{U}|\psi_{i}\rangle$. Such a process obviously exactly simulates the sampling from the density matrix $\hat{\rho}=\sum_{i}p_{i}\hat{U}|\psi_{i}\rangle\langle\psi_{i}|\hat{U}^{\dagger}$. A subtlety with this approach is that, in principle, many different decompositions are consistent with the given density matrix, and different decompositions may incur different computational costs. An obvious optimal choice for an efficient MPS simulation is a decomposition such that each of the pure-state components $|\psi_{i}\rangle$ has a small entanglement throughout the circuit.
For this particular setup of lossy boson sampling, one naive but natural choice of the decomposition of $\hat{\sigma}$ is a convex sum of $|0\rangle\langle 0|$ and $|1\rangle\langle 1|$ with probabilities $1-\eta$ and $\eta$, respectively. However, one may easily notice a drawback of this decomposition, which leads to a significant computational cost [31]. More specifically, since the transmission rate is $\eta$, the sampled pure state input contains an average $\eta N$ number of input $|1\rangle\langle 1|$. To simulate this sampling problem in polynomial time, $\eta N=O(\log N)$ is required because the complexity of boson sampling is exponential to the number of single photon inputs. Therefore, the simulable range of $\eta$ is too limited compared to the known simulable range $\eta=O(1/\sqrt{N})$ [25].
We now introduce another decomposition of the lossy single photon state $\hat{\sigma}$ and prove that it provides a wider range of simulability, which is given as
$\hat{\sigma}=\dfrac{1}{2}|\psi_{+}\rangle\langle\psi_{+}|+\dfrac{1}{2}|\psi_{-}\rangle\langle\psi_{-}|,$ (17)
where we defined
$|\psi_{\pm}\rangle\equiv\sqrt{1-\eta}|0\rangle\pm\sqrt{\eta}|1\rangle.$ (18)
The decomposition indicates that the lossy single photon state input $\hat{\sigma}$ is physically equivalent to generating two pure state inputs $|\psi_{+}\rangle$ and $|\psi_{-}\rangle$ with equal probability. Hence, the total input state can be rewritten as
$\hat{\rho}_{\text{in}}=\sum_{s_{1},\dots,s_{N}\in\{-,+\}^{N}}\frac{1}{2^{N}}|\psi_{\text{in}}(s_{1},\dots,s_{N})\rangle\langle\psi_{\text{in}}(s_{1},\dots,s_{N})|,$ (19)
where
$|\psi_{\text{in}}(s_{1},\dots,s_{N})\rangle\equiv\left[\bigotimes_{j=1}^{N}|\psi_{s_{j}}\rangle\right]\otimes|0\rangle^{\otimes(M-N)}.$ (20)
Therefore, simulating $\hat{\rho}_{\text{in}}$ is equivalent to sampling $(s_{1},\dots,s_{N})\in\{-,+\}^{N}$ from a uniform probability and then simulating the circuit with the input $|\psi_{\text{in}}(s_{1},\dots,s_{N})\rangle$. To demonstrate that such a decomposition indeed results in low computational complexity for simulation, the following subsection analyzes the computational cost of our decomposition using the ERE.
Before we analyze the entanglement, note that $|\psi_{-}\rangle$ is equivalent to $e^{i\pi\hat{n}}|\psi_{+}\rangle$, where $e^{i\pi\hat{n}}$ is a linear-optical circuit. Hence, for the simplicity of the analysis, we fix the sampled lossy input state to be $|\psi_{+}\rangle^{\otimes N}\otimes|0\rangle^{\otimes(M-N)}$, without loss of generality for our analysis, since the Haar-random unitary ensemble is invariant under the unitary multiplication (corresponding to the phase shifter $e^{i\pi\hat{n}}$).
Figure 4: Simulation of lossy boson sampling using pure-state decomposition. (a) The Fock input $|1\rangle^{\otimes N}|0\rangle^{\otimes(M-N)}$ goes through the loss channel, yielding the product mixed state $\hat{\sigma}^{\otimes N}\otimes|0\rangle\langle 0|^{\otimes(M-N)}$, where $\hat{\sigma}$ is the lossy single-photon state. (b) For sampling, each $\hat{\sigma}$ is decomposed into $|\psi_{+}\rangle$ and $|\psi_{-}\rangle$, drawn with equal probability.
IV.3 Entanglement Rényi entropy of the output state
The main reason we choose the above decomposition is that each of the pure state components in the decomposition can constitute only a small amount of entanglement from a linear-optical circuit when a loss rate is large enough. To see this, let us now derive the entanglement of the output state of each $|\psi_{\text{in}}\rangle$ after a linear-optical circuit $\hat{U}$ and the computational complexity for the MPS simulation depending on the transmission rate $\eta$.
Let us first analyze the output state of a linear-optical circuit for input $|\psi_{\text{in}}\rangle$. After the linear-optical circuit $\hat{U}$, the state transforms as follows:
$|\psi_{\text{in}}\rangle$ $=\left[\bigotimes\limits_{j=1}^{N}\left(\sqrt{1-\eta}|0\rangle+\sqrt{\eta}|1\rangle\right)\right]\otimes|0\rangle^{\otimes(M-N)}$ (21)
$=\left[\prod\limits_{j=1}^{N}\left(\sqrt{1-\eta}+\sqrt{\eta}\hat{a}_{j}^{\dagger}\right)\right]|0\rangle^{\otimes M}$ (22)
$\rightarrow\left[\prod\limits_{j=1}^{N}\left(\sqrt{1-\eta}+\sqrt{\eta}\hat{b}_{j}^{\dagger}\right)\right]|0\rangle^{\otimes M}\equiv|\psi_{\text{out}}\rangle,$ (23)
where the linear-optical circuit $\hat{U}$ transforms the creation operator $\hat{a}_{j}^{\dagger}$ into $\hat{b}_{j}^{\dagger}\equiv\hat{U}^{\dagger}\hat{a}_{j}^{\dagger}\hat{U}$. Let us now investigate the entanglement of the state between a bipartition $1,\dots,l]$ and $(l+1),\dots,M]$. To do so, we trace out one of the two subsystems and derive the ERE of the reduced density matrix. Note that we can rewrite the bosonic creation operator $\hat{b}_{j}^{\dagger}$ as a sum of operators on each partition
$\hat{b}_{j}^{\dagger}=\sum\limits_{k=1}^{M}U_{jk}^{*}\hat{a}_{k}^{\dagger}\equiv\cos\theta_{j}\hat{B}_{\text{u},j}^{\dagger}+\sin\theta_{j}\hat{B}_{\text{d},j}^{\dagger},$ (24)
where we defined
$\cos\theta_{j}\hat{B}_{\text{u},j}^{\dagger}=\sum\limits_{k=1}^{l}U_{jk}^{*}\hat{a}_{k}^{\dagger},\penalty 10000\ \penalty 10000\ \sin\theta_{j}\hat{B}_{\text{d},j}^{\dagger}=\sum\limits_{k=l+1}^{M}U_{jk}^{*}\hat{a}_{k}^{\dagger}.$ (25)
Here, u means the upper spatial bipartition $1,\dots,l]$, and d means the other partition.
In this setting, $\cos^{2}\theta_{j}$ can be interpreted as the probability that a photon in output mode $j$ is located within the first subsystem after the linear-optical circuit $U$ (i.e., within modes $1,\dots,l]$), and $\sin^{2}\theta_{j}$ is the corresponding probability for the second subsystem. By definition, the normalization is given as
$\cos^{2}\theta_{j}=\sum\limits_{k=1}^{l}|U_{jk}|^{2},\penalty 10000\ \penalty 10000\ \penalty 10000\ \sin^{2}\theta_{j}=\sum\limits_{k=l+1}^{M}|U_{jk}|^{2}.$ (26)
Under the assumption that $M=\omega(N^{2})$, a typical assumption for the hardness result of boson sampling, the creation operators $\hat{B}_{\text{u},j}^{\dagger}$ and $\hat{B}_{\text{d},j}^{\dagger}$ follow the canonical commutation relation approximately,
$[\hat{B}_{\text{u},j},\hat{B}_{\text{u},k}^{\dagger}]\approx\delta_{jk},\penalty 10000\ \penalty 10000\ \penalty 10000\ [\hat{B}_{\text{d},j},\hat{B}_{\text{d},k}^{\dagger}]\approx\delta_{jk},$
$[\hat{B}_{\text{u},j},\hat{B}_{\text{d},k}^{\dagger}]=0,\penalty 10000\ \penalty 10000\ \penalty 10000\ [\hat{B}_{\text{u},j}^{\dagger},\hat{B}_{\text{d},k}]=0.$ (27)
We will discuss this point later.
Using those tools, we now evaluate the ERE between the bipartite subsystems of the same size, where the ERE is maximum [48, 49]; therefore, $l$ is chosen to be $\lfloor N/2\rfloor$. To do that, we now compute the eigenvalues of the reduced density matrix on the subsystem u (or equivalently the subsystem d).
Substituting the splitted operators transforms the output state $|\psi_{\text{out}}\rangle$ as
$\bigotimes\limits_{j=1}^{N}\left[\left(\sqrt{1-\eta}+\sqrt{\eta}\left(\cos\theta_{j}\hat{B}_{\text{u},j}^{\dagger}+\sin\theta_{j}\hat{B}_{\text{d},j}^{\dagger}\right)\right)|0\rangle\right]$ (28)
$=\bigotimes\limits_{j=1}^{N}\left[\sqrt{1-\eta}|00\rangle_{j}+\sqrt{\eta}\cos\theta_{j}|10\rangle_{j}+\sqrt{\eta}\sin\theta_{j}|01\rangle_{j}\right].$ (29)
For $j>N$, the modes remain in the vacuum state $|00\rangle_{j}$, which do not affect the entanglement and are therefore omitted. For brevity, we also omit the labels of subsystem u and d when it is clear from the context; for example, we write $\hat{B}_{\text{u},j}^{\dagger}|0\rangle=|1\rangle_{\text{u},j}|0\rangle_{\text{d},j}$ simply as $|10\rangle_{j}$, where the first and second entry refers to subsystems u and d, respectively.
To derive the ERE between the two bipartitions, we find the eigenvalues of the reduced density matrix of $|\psi_{\text{out}}\rangle\langle\psi_{\text{out}}|$. After tracing out the subsystem d, the reduced density matrix on the subsystem u can be represented as the following matrix:
$\bigotimes_{j=1}^{N}\begin{pmatrix}1-\eta\cos^{2}\theta_{j}&\sqrt{\eta(1-\eta)}\cos\theta_{j}\\ \sqrt{\eta(1-\eta)}\cos\theta_{j}&\eta\cos^{2}\theta_{j}\end{pmatrix}$, (30)
where we have omitted the trivial tensor product $|0\rangle\langle 0|^{\otimes(M-N)}$ as above. Therefore, the eigenvalues of the reduced density matrix are given as $\prod_{j=1}^{N}c_{jk_{j}}$, where $\vec{k}\in\{0,1\}^{N}$ and
$c_{j0}=\dfrac{1}{4}\left(2+\sqrt{2\eta^{2}\cos(4\theta_{j})-2\eta^{2}+4}\right),$ (31)
$c_{j1}=\dfrac{1}{4}\left(2-\sqrt{2\eta^{2}\cos(4\theta_{j})-2\eta^{2}+4}\right).$ (32)
To determine the classically simulable range of $\eta$, we use the criterion that the ERE scales as $O(\log N$) for $0<\alpha<1$. One may easily see that the ERE is maximized when $\theta_{j}=\pi/4$, where the deviation from $1/2$ of the two eigenvalues of each matrix is minimal. In this case, the ERE is given by
$S_{\alpha}^{l}(|\psi_{\text{out}}\rangle)$
[…]
which is bounded by a simpler expression with respect to $\eta$:
[…]
$\leq\dfrac{N}{1-\alpha}\left(\dfrac{1}{2}\right)^{\alpha}\left(1-\sqrt{1-\eta^{2}}\right)^{\alpha}$ (35)
$\leq\dfrac{N\eta^{2\alpha}}{1-\alpha}.$ (36)
From the first line to the second line, we applied the inequality $\log(1+x)\leq x$, which holds when $x\geq 0$. Since the inequality $1-\sqrt{1-x^{2}}\leq x^{2}$ holds when $1\leq x\leq 1$, it can be applied to the second line to derive the third line. Recalling that the state is efficiently simulable when the above upper bound of the ERE is $O(\log N)$ for $0<\alpha<1$, we conclude that when
$\eta$ $=O\left(\left(\dfrac{\log N}{N}\right)^{1/2\alpha}\right),$ (37)
the quantum state after linear-optical circuits is efficiently described and simulated using MPS, where $\alpha$ can be chosen arbitrarily close to $1$.
On the other hand, we can also obtain the lower bound of $\eta$, which makes MPS inapproximable:
$S_{\alpha}^{l}(|\psi_{\text{out}})\rangle$ $\geq\dfrac{N}{1-\alpha}\log\left[\left(1-\dfrac{\eta^{2}}{2}\right)^{\alpha}\right]$ (38)
$\geq\dfrac{\alpha N\eta^{2}}{2(\alpha-1)}.$ (39)
At the first line, $1+\sqrt{x}\geq 1+x$ for $0\leq x\leq 1$ is used. From the first line to the second line, $\log(1-x)\geq-x$ when $x\leq 0.863$ is used, which is always valid since $\eta\leq 1$. Hence, when
$\eta=\Omega\left(\frac{1}{N^{1/2-\kappa/2}}\right),$ (40)
where $\kappa>0$ is a constant, the ERE is lower bounded by $\Omega(N^{\kappa})$. Hence, by choosing $\kappa\to 0$, i.e., $\eta$ scales faster than $1/\sqrt{N}$, the corresponding lossy boson sampling is hard to simulate using our MPS method. Therefore, the approximability achieved by our MPS method exhibits a transition at $\eta=\Theta(1/\sqrt{N})$, which is consistent with the GBS case [33].
To clearly identify the transition point, we numerically illustrate this behavior using the von Neumann entanglement entropy, varying the coefficient of transmission rate over $1/\sqrt{N}$. We can see that between the coefficients $4$ and $5$, the transition from approximability to inapproximability of the state occurs. One may see that when the transmission rate is smaller than $\eta=4/\sqrt{N}$, the coefficient of the $\log N$ term in the von Neumann entanglement entropy decreases as we increase the system size. In contrast, the transmission rate is larger than $\eta=5/\sqrt{N}$, the coefficient of the $\log N$ term in the von Neumann entanglement entropy increases as the system size grows. Hence, it indicates that the transition occurs between $\eta=4/\sqrt{N}$ and $5/\sqrt{N}$.
Figure 5: Entanglement (von Neumann) entropy of the output state of lossy boson sampling in the worst case, $\theta=\pi/4$. A complexity transition, from efficient to inefficient, occurs between coefficient values 4 and 5.
We now compare our results with those of previous studies [25, 24]. In the previous work, classical algorithms are proposed whose classically simulable range of $\eta$ is $\eta=O(1/\sqrt{N})$. The algorithms employed a method that approximates the lossy input state as a state that is easy to classically simulate, such as the closest thermal state or the closest separable state. An obvious crucial limitation of this approach is that the closest thermal state or the closest separable state to a given lossy input state remains fixed once the parameters of the lossy boson sampling are determined. That means that when the system is not sufficiently lossy, no thermal state or separable state can approximate the lossy input state with adequate accuracy. Furthermore, the simulation accuracy cannot be improved by allocating additional computational resources because the closest easy-to-simulate state is fixed.
Compared to that, although our MPS algorithm exhibits a similar scaling in an asymptotic regime as $\eta<O(1/\sqrt{N})$, i.e., our MPS method might not provide an advantage over the previously known classical algorithm asymptotically, it shows a wider simulable range than the previous work in a non-asymptotic regime, which is more practically relevant, because even if the loss rate is not high enough, our algorithm always runs and can achieve a target accuracy by using more computational resources, i.e., by increasing the bond dimension. We emphasize that this improvement significantly enhances simulation performance in simulating lossy GBS [26, 33, 50]. To be more specific, Ref. [26] shows that the closest nonclassical state renders a similar classically simulable range of $\eta$, $\eta=O(1/\sqrt{N})$, which, however, was insufficient to simulate the experimental GBS. Nonetheless, Ref. [33] developed a classical algorithm based on MPS, just like ours, and could simulate the state-of-the-art experimental GBS, which clearly indicates that the MPS method significantly outperforms other methods in practice even though the asymptotic scaling is similar.
We now compare this with another previous result based on MPO, which numerically finds a similar classically simulable range $\eta=O(1/\sqrt{N})$ [27]. Although the previous result yields the same strength in the sense that the accuracy can be improved by adding more computational resources, our work demonstrates a significant enhancement of the numerical simulability range. More specifically, the computational time cost of MPS and MPO simulations in a circuit of depth $D$ is $T=O(MDd^{4}\chi^{3})$ and $T=O(MDd^{8}\chi^{3})$, respectively. In terms of memory requirement, the memory cost of MPS is $O(\chi^{2}dM)$, while that of MPO is $O(\chi^{2}d^{2}M)$. In the MPO case, the local Hilbert space dimension $d$ is the square of that in the MPS case. This can lead to a higher computational cost when $d=N+1$ and $N$ is sufficiently large, as is the case in the boson sampling scenario. More crucially, as mentioned earlier, classical correlation can contribute to the increase in the bond dimension of MPO, whereas it cannot for MPS. Therefore, the bond dimensions between the two methods also differ between MPS and MPO. For example, the MPO method requires a bond dimension of the order of $10^{4}$, even for an input photon of less than $10$, with a loss rate of $0.7$. Our MPS method, on the other hand, requires only a bond dimension of the order of $10$ (note that the definitions of approximation errors differ, but a significant difference in bond dimension is evident).
Let us finally discuss the canonical commutation relation of $\hat{B}_{\text{u}}$ and $\hat{B}_{\text{d}}$. For our analytic results, we assumed that they satisfy the exact canonical commutation relation as in Eq. (27).
$\left[\hat{B}_{\text{u},j},\hat{B}_{\text{u},k}^{\dagger}\right]$ $=\left[\dfrac{\sum\limits_{m=1}^{l}U_{jm}\hat{a}_{m}}{\cos\theta_{j}},\dfrac{\sum\limits_{n=1}^{l}U_{kn}^{*}\hat{a}_{n}^{\dagger}}{\cos\theta_{k}}\right]$ (41)
$=\dfrac{\sum\limits_{m=1}^{l}U_{jm}U_{km}^{*}}{\sqrt{\sum\limits_{m=1}^{l}|U_{jm}|^{2}}\sqrt{\sum\limits_{n=1}^{l}|U_{kn}|^{2}}}.$ (42)
For $j=k$, this is equal to 1 for all unitary. However, for $j\neq k$, this is not exactly 0, but we can show that it is approximately 0 in the Haar random case. See Appendix B for details. More specifically, we show that $\left[\hat{B}_{\text{u},j},\hat{B}_{\text{u},k}^{\dagger}\right]=O(1/\sqrt{M})$ with high probability. Thus, our approximation is valid when $M$ is sufficiently large.
We also justify this commutation relation without the Haar-random assumption. It has been known that in the lossless case, the linear-optical circuit that renders the maximum entanglement entropy is the one that transforms the initial bosonic creation operators as [33, 35]
$\hat{a}_{i}^{\dagger}\to\frac{1}{\sqrt{2}}(\hat{a}_{i}^{\dagger}+\hat{a}_{i+M/2}^{\dagger}),$ (43)
for $1\leq i\leq N$, where $M$ is assumed to be even for simplicity. In this case, the assumed commutation relations exactly hold. Although we do not prove that this also holds in the lossy case, we conjecture that this is true.
IV.4 Numerical analysis of the computational cost
To understand how large parameter regimes our classical algorithm can simulate in practice, we now directly estimate the bond dimension for simulating the boson sampling algorithm with the photon number $N$ and transmission rate $\eta$, with a fixed MPS approximation error $\epsilon$ of $0.01$. Our MPS approximation error is defined by the sum of the discarded eigenvalues when the system is bipartitioned into two halves. Before we present our results, we emphasize that for certain regimes, other existing methods, such as the Clifford-Clifford algorithm, may outperform our method [23, 39]. Hence, one should interpret our results as a boundary set by our MPS method, rather than the boundary for the classically simulable regime for all possible methods.
In this numerical analysis, we consider three different families of circuits: (a) the worst-case construction, (b) the Haar-random unitary, and (c) an experimentally motivated structure. The purpose of this comparison is to examine how the required bond dimension varies across different choices of circuit unitaries. Here, the worst-case construction has $\theta_{j}=\pi/4$ for all $j$ to produce the largest amount of entanglement between the bipartitions we consider, and the experimentally motivated structure follows the same structure implemented in Ref. [12]. For (c), although the experiment in the reference implements GBS, we constructed a circuit unitary from this setup. Since the number of modes of this unitary is restricted as $M=2k^{2}$ for an integer $k$, for a given photon number $N$, we set $k$ as the nearest integer to $N/\sqrt{2}$ so that $M=2k^{2}$ is the closest value to $N^{2}$. However, we inject $\lceil N/2\rceil$ photons into the middle of the modes of the first half, and the remaining $\lfloor N/2\rfloor$ photons into the middle of the modes of the second half. This is because the whole unitary in this case is constructed by combining two unitaries of half the size; inputting the same number of single photons in both bipartitions and inputting single photons in the middle can help maximize the entanglement.
We present the numerical results in Fig. 6. First of all, Fig. 6(a) presents the required bond dimension in the worst-case circuit. Fig. 6(b) is when the unitary of the boson sampling circuit is drawn from a Haar random. The required resource in this result is comparable to the worst-case scenario, indicating that highly entangled configurations are typical in Haar random cases. One can see that if the loss rate is approximately $50\%$, which is the case for GBS experiments in Refs. [13, 15, 17], and the input photon number is $40$, the required bond dimension is $10^{7}$, which is already quite large to implement in practice.
In contrast to the previous cases, Fig. 6(c) shows a significant decrease in the required bond dimension. For example, when the transmission rate is $0.5$ and the input photon number is $40$, the required bond dimension in Fig. 6(a) and (b) reaches $10^{7}$, indicating that Haar-random unitaries approximately realize the worst-case scaling. In contrast, Fig. 6(c) requires less than $10^{6}$ bond dimensions under the same conditions, which is an order of magnitude smaller. Since the asymptotic simulation memory cost and time cost scale with the square and cube of the bond dimension, respectively, this $10$ times reduction in Fig. 6(c) translates into nearly $100$ times fewer memory, and $1000$ times fewer calculation time.
Figure 6: Required bond dimension and memory to achieve an MPS approximation error $\epsilon=0.01$. Panels (a)–(c) show the required bond dimension, and panels (d)–(f) show the required memory. (a) Required bond dimension where the unitary is in the worst-case scenario. This implies the highest cost for a specific input photon number and photon loss. (b) Required bond dimension when the Haar random unitary is a unitary corresponding to the circuit in boson sampling. (c) Required bond dimension when a circuit unitary is the same structure as the GBS experiment of USTC [12]. (d) Required memory where the unitary is in the worst-case scenario. (e) Required memory when the Haar random unitary is a unitary corresponding to the circuit in boson sampling. (f) Required memory when a circuit unitary is the same structure as the GBS experiment of USTC [12].
According to the required bond dimension, the memory complexity of the circuit can be estimated as shown in Fig. 6(d)–(f) [33]. Assuming that each complex number requires 8 bytes, the memory usage can be estimated as the total number of elements multiplied by 8 bytes:
$(\text{8 bytes})\times(\text{bond dimension})^{2}\times(\text{number of modes})$
$\times(\text{local Hilbert space dimension}).$ (44)
Here, when the input state of boson sampling is $N$ single photons, the local Hilbert space dimension can be considered as $N+1$ in the general case.
IV.5 Multiphoton Fock boson sampling
We now generalize the MPS method for single-photon boson sampling to multiphoton Fock boson sampling. Suppose that for the first $N$ modes, the input is each $n$ photon Fock state, and the MNM-N remaining modes are initialized in the vacuum state. The linear optical structure and the detectors are the same as in the previous boson sampling case. After bringing the noise to the front of the circuit as in Fig. 1 and passing the input state through the photon loss channel with total transmission rate $\eta$, the lossy input state becomes (see Fig. 4)
$\hat{\rho}_{\text{in}}=\hat{\sigma}^{\otimes N}\otimes|0\rangle\langle 0|^{\otimes(M-N)},$ (45)
where the lossy Fock state is written as
$\hat{\sigma}=\sum_{k=0}^{n}\dfrac{n!}{k!(n-k)!}(1-\eta)^{n-k}\eta^{k}|k\rangle\langle k|.$ (46)
To utilize the MPS formalism, as before, we first find an appropriate pure-state decomposition and sample a pure state from the given mixed state. Our choice of the decomposition is given by the states
$|\psi(\phi)\rangle$ $=\sum_{k=0}^{n}\sqrt{\binom{n}{k}\eta^{k}(1-\eta)^{n-k}}e^{i\phi\hat{a}^{\dagger}\hat{a}}|k\rangle,$ (47)
with $\phi$ sampled uniformly from $0,2\pi)$. One can easily verify that $\frac{1}{2\pi}\int_{0}^{2\pi}d\phi\,|\psi(\phi)\rangle\langle\psi(\phi)|=\hat{\sigma}$. Hence, with this decomposition, to simulate the state $\hat{\sigma}$, we can instead sample $|\psi(\phi)\rangle$ following the uniform distribution over $0,2\pi)$. Thus, the sampled state for all modes then takes the form
$\left[\bigotimes_{j=1}^{N}|\psi(\phi_{j})\rangle\right]\otimes|0\rangle^{\otimes(M-N)},$ (48)
for sampled values of $\{\phi_{j}\}_{j=1}^{N}$. As in the single-photon boson sampling case, since a single layer of phase shifters can align all phases $\phi_{j}$ to the same value, and this layer can be absorbed into the linear-optical circuit, we may take $\phi_{j}=0$ for all $j$ without loss of generality. Accordingly, the sampled input state is written as
$|\psi(0)\rangle^{\otimes N}\otimes|0\rangle^{\otimes(M-N)}.$ (49)
After the transformation of the creation operator as a lossy single-photon case, $\hat{a}_{j}^{\dagger}\rightarrow\hat{b}_{j}^{\dagger}=\cos\theta_{j}\hat{B}_{\text{u},j}^{\dagger}+\sin\theta_{j}\hat{B}_{\text{d},j}^{\dagger}$, we have the output state as follows:
$|\psi(0)\rangle$ $\rightarrow\sum\limits_{k=0}^{n}\sqrt{\binom{n}{k}\eta^{k}(1-\eta)^{n-k}}\dfrac{(\hat{b}^{\dagger})^{k}}{\sqrt{k!}}|0\rangle$ (50)
$=\sum\limits_{k=0}^{n}\sqrt{{n\choose k}\eta^{k}(1-\eta)^{n-k}}$ (51)
$\quad\times\sum\limits_{l=0}^{k}\sqrt{{k\choose l}}\cos^{l}\theta\sin^{k-l}\theta|l,k-l\rangle.$ (52)
After tracing out the down part, the eigenvalues of the reduced state can be calculated to evaluate the ERE, which serves as an indicator of classical simulability. Since the closed-form of the ERE is difficult to find, unlike the single photon cases, we calculate it numerically and present the behavior in Fig. 7. Figure 7 presents the ERE relative to $\log N$ in terms of the number of modes $N$ with the $n$-photon Fock state for each mode with different scaling of transmission rate $\eta$. Note that when the slope of the graph decreases, the corresponding ERE scales as $O(\log N)$, which means that MPS efficiently describes the state. Thus, by observing Fig. 7(a), we can conclude that all the different Fock state inputs are efficiently simulable by MPS when $\eta=1/\sqrt{N}$. In Fig. 7(b), when $\alpha<1$, when the transmission rate $\eta$ scales equal or smaller than $1/\sqrt{N}$, the state is efficiently simulable by MPS, which is consistent with the result of the single photon boson sampling case. On the other hand, when the circuit has less loss, $\eta=1/N^{1/3}$, its ERE scales faster than $O(\log N)$, which means that it is not determined whether MPS efficiently describes it. The inapproximability of this $\eta=1/N^{1/3}$ case can be determined by analyzing the $\alpha>1$ case, as in Fig. 7(c). For $\alpha>1$, when the entropy scales as $\Omega(N^{\kappa})$ for $\kappa>0$, the state is not simulable. Here, the graph scales linearly in a log–log plot, which means there exists a proper $\kappa$ for the Rényi entropy scaling. Thus, this case is inapproximable by MPS. Hence, we observe a similar behavior for general Fock-state boson sampling to the single-photon boson sampling case.
Figure 7: ERE with respect to input photon number $N$, for different photon numbers in one mode $n$ and transmission rates $\eta$. (a) The coefficient of the $\log N$ term in the ERE decreases as $n$ increases. (b) For $\eta=1/\sqrt{N}$, the ERE scales as $O(\log N)$, which is efficiently described by MPS; however, when $\eta=1/(N^{1/3})$, the ERE does not scale as $O(\log N)$. It is unclear whether an MPS description remains efficient. (c) The regime that was unclear in (b) is shown to be inefficient for an MPS by examining the scaling when $\alpha>1$.
IV.6 Cat-state boson sampling
Another generalization of the boson sampling problem exploits a cat state input, where the even and odd cat states are defined as
$|\text{cat}_{e}\rangle$ $\equiv\frac{1}{\sqrt{2\left(1+e^{-2|\gamma|^{2}}\right)}}(|\gamma\rangle+|-\gamma\rangle),$ (53)
$|\text{cat}_{o}\rangle$ $\equiv\frac{1}{\sqrt{2\left(1-e^{-2|\gamma|^{2}}\right)}}(|\gamma\rangle-|-\gamma\rangle),$ (54)
respectively. Note that as the amplitude of the cat state $\gamma\rightarrow 0$, the odd (even) cat state approaches $|1\rangle$ ($|0\rangle$), thereby recovering single-photon inputs. Moreover, cat state sampling is also known to be hard to classically simulate [51]. The circuit has the same structure as before, except that the input of the first $N$ mode is an odd cat state. The photon loss can be gathered in front of the circuit as in Fig. 1, and the cat state input goes through the total photon loss channel and transforms to the following state,
$\hat{\rho}_{\text{in}}=\hat{\sigma}^{\otimes N}\otimes|0\rangle\langle 0|^{\otimes(M-N)},$ (55)
where the noisy cat state becomes
$\hat{\sigma}$ $\equiv\frac{1}{C^{2}}\left[|\gamma\sqrt{\eta}\rangle\langle\gamma\sqrt{\eta}|-e^{-2|\gamma|^{2}(1-\eta)}(|\gamma\sqrt{\eta}\rangle\langle-\gamma\sqrt{\eta}|\right$.
$+\left.|-\gamma\sqrt{\eta}\rangle\langle\gamma\sqrt{\eta}|)+|-\gamma\sqrt{\eta}\rangle\langle-\gamma\sqrt{\eta}|\right],$ (56)
where $C\equiv\sqrt{2(1-e^{-2|\gamma|^{2}})}$ is defined for the normalization. As before, we now find a pure-state decomposition of the lossy input state $\hat{\sigma}$. Our choice of decomposition is as follows:
$\hat{\sigma}=\dfrac{1}{2}|\psi_{1}\rangle\langle\psi_{1}|+\dfrac{1}{2}|\psi_{2}\rangle\langle\psi_{2}|,$ (57)
where we defined
$|\psi_{1}\rangle$ $\equiv A|\gamma\sqrt{\eta}\rangle-B|-\gamma\sqrt{\eta}\rangle,$ (58)
$|\psi_{2}\rangle$ $\equiv-B|\gamma\sqrt{\eta}\rangle+A|-\gamma\sqrt{\eta}\rangle,$ (59)
$A\equiv$ $\dfrac{\sqrt{1+k}+\sqrt{1-k}}{\sqrt{2}C},B\equiv\dfrac{\sqrt{1+k}-\sqrt{1-k}}{\sqrt{2}C},$ (60)
with $k=e^{-2|\gamma|^{2}(1-\eta)}$.
Following the method illustrated in Fig. 4, the mixed state is equivalent to generating pure states $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$ with each probability $1/2$. As in the previous sections, in the following calculations, the sampled input state $|\psi_{1}\rangle^{\otimes N}\otimes|0\rangle^{\otimes(M-N)}$ is used without loss of generality because $|\psi_{2}\rangle=e^{i\pi\hat{n}}|\psi_{1}\rangle$ as in the single-photon boson sampling case.
We can calculate the output state after the circuit. Using the same transformation of creation operator, $\hat{a}_{j}^{\dagger}\rightarrow\hat{b}_{j}^{\dagger}=\cos\theta_{j}\hat{B}_{\text{u},j}^{\dagger}+\sin\theta_{j}\hat{B}_{\text{d},j}^{\dagger}$ as the preceding sections, the input state evolves as
$|\psi_{1}\rangle$ $=A|\gamma\sqrt{\eta}\rangle-B|-\gamma\sqrt{\eta}\rangle$ (61)
$\rightarrow A|\gamma\sqrt{\eta}\cos\theta,\gamma\sqrt{\eta}\sin\theta\rangle$ (62)
$\quad-B|-\gamma\sqrt{\eta}\cos\theta,-\gamma\sqrt{\eta}\sin\theta\rangle.$ (63)
Let us define $\delta\equiv\gamma\sqrt{\eta}\cos\theta$ for simplicity. To analyze the ERE of the output state subsystem, the subsystem d is traced out, then the reduced density matrix is written as
$\hat{\rho}_{\text{u}}$ $=\text{tr}_{\text{d}}(|\psi_{1}\rangle\langle\psi_{1}|)$ (64)
$=A^{2}|\delta\rangle\langle\delta|-ABe^{-2|\gamma|^{2}\eta\sin^{2}\theta}(|\delta\rangle\langle-\delta|$ (65)
$\quad+|-\delta\rangle\langle\delta|)+B^{2}|-\delta\rangle\langle-\delta|,$ (66)
Diagonalizing this in the orthogonal even/odd cat basis yields the eigenvalues
$\lambda_{\pm}\equiv\dfrac{1}{2}\left(1\pm\sqrt{1-\zeta}\right),$ (67)
where
$\zeta\equiv\dfrac{e^{-4|\gamma|^{2}+4|\gamma|^{2}\eta}(1-e^{-4|\gamma|^{2}\eta\cos^{2}\theta})(1-e^{-4|\gamma|^{2}\eta\sin^{2}\theta})}{(1-e^{-2|\gamma|^{2}})^{2}}.$ (68)
Thus, ERE of the state $|\psi_{\text{out}}\rangle$ is given by
[…]
Recall that when the ERE scales $O(\log N)$ for $0<\alpha<1$, the state is classically simulable in polynomial time.
Here, the ERE reaches its maximum value when $\theta=\pi/4$. With this, we can obtain the upper bound of the ERE and determine the asymptotically simulable range. Applying this condition to bound $S_{\alpha}^{l}(|\psi_{\text{out}}\rangle)$,
$S_{\alpha}^{l}(|\psi_{\text{out}}\rangle)\leq\dfrac{N}{1-\alpha}\left(\dfrac{1}{2}\right)^{\alpha}\zeta^{\alpha}$ (70)
$=\dfrac{N}{1-\alpha}\dfrac{1}{2}\left(\dfrac{1}{2}\right)^{\alpha}\left[\dfrac{(1-e^{-2|\gamma|^{2}\eta})^{2}(1-e^{-4|\gamma|^{2}+4|\gamma|^{2}\eta})}{(1-e^{-2\gamma^{2}})^{2}}\right]^{\alpha}$ (71)
$\leq\dfrac{N}{1-\alpha}\left(\dfrac{1}{2}\right)^{\alpha+1}\left[\dfrac{(2|\gamma|^{2}\eta)^{2}(4|\gamma|^{2}-4|\gamma|^{2}\eta)}{(1-e^{-2\gamma^{2}})^{2}}\right]^{\alpha}$ (72)
At the first line, the same inequalities as boson sampling, Eq. (34), Eq. (35), and Eq. (36) are used. From the second line to the third line, $1-e^{-x}\leq x$ for $x\geq 0$ is used. When this ERE scales as $O(\log N)$, the MPS of the state is classically efficiently simulable, namely,
$\eta$ $=O\left(\left(\dfrac{\log N}{N}\right)^{1/2\alpha}\right),$ (73)
which is consistent with the standard boson sampling case. For even cat state input, the eigenvalues are the same except that the denominator is $(1+e^{-2|\gamma|^{2}})^{2}$, and the classically efficiently simulable $\eta$ is asymptotically the same.
V Noisy IQP sampling
V.1 Decomposition of noisy input state
We now demonstrate that our MPS-based classical simulation strategy developed for boson sampling can be extended to IQP sampling. The central idea is similar to that of boson sampling, in that we find a decomposition of a noisy input state into pure states, such that the entanglement of each pure state after the IQP circuit is low. As introduced in Sec. II.2, we consider an IQP circuit with $n$ input qubits prepared in the state $|+\rangle$. We first focus on the dephasing noise case and then extend our analysis to the depolarizing noise. When the circuit has depth $d$ and each layer suffers dephasing noise with noise rate $p$, all dephasing noise channels $\mathcal{N}_{0,0,p}$ can be commuted to the front of the diagonal gates, resulting in a single layer of a dephasing channel $\mathcal{N}_{0,0,p_{d}}$. As illustrated in Fig. 2, the noisy circuit can therefore be decomposed into an overall dephasing noise channel acting on the input, followed by an ideal IQP circuit without dephasing.
After the input state $|+\rangle^{\otimes n}$ goes through the noise channel $\mathcal{N}_{0,0,p_{d}}$, the input state transforms to noisy states
$\hat{\rho}_{\text{in}}=\hat{\tau}^{\otimes n},$ (74)
where we defined the noisy $|+\rangle$ state under dephasing channel as
$\hat{\tau}\equiv\mathcal{N}_{0,0,p_{d}}(|+\rangle\langle+|)=(1-p_{d})|+\rangle\langle+|+p_{d}|-\rangle\langle-|.$ (75)
As in boson sampling, the noisy input state $\hat{\tau}$ can be represented as a probabilistic mixture of pure states, and this sampling property enables us to represent the resulting pure state by MPS and thus simulate the noisy IQP circuit. Again, for a better performance, we find an appropriate decomposition of the mixed state such that each pure state from the decomposition renders a small enough entanglement at the end of the circuit, which admits an efficient classical description. A naive approach is to sample this state by $|+\rangle$ with probability $1-p_{d}$, and into $|$-$\rangle$ with probability $p_{d}$. However, this does not reduce the complexity of the classical simulation because $|$-$\rangle$ is a $|+\rangle$ state with a $Z$ gate applied, and thus we eventually have to simulate the noiseless IQP circuit in any case. Therefore, this trivial decomposition does not reduce the simulation complexity.
Instead, one may easily see that the noisy input state can be decomposed as
$\hat{\tau}=\frac{1}{2}|\phi_{+}\rangle\langle\phi_{+}|+\frac{1}{2}|\phi_{-}\rangle\langle\phi_{-}|,$ (76)
where we defined
$|\phi_{+}\rangle\equiv\sqrt{1-p_{d}}|+\rangle+\sqrt{p_{d}}|-\rangle=q_{0}|0\rangle+q_{1}|1\rangle,$ (77)
$|\phi_{-}\rangle\equiv\sqrt{1-p_{d}}|+\rangle-\sqrt{p_{d}}|-\rangle=q_{1}|0\rangle+q_{0}|1\rangle,$ (78)
with
$q_{0}\equiv\dfrac{\sqrt{1-p_{d}}+\sqrt{p_{d}}}{\sqrt{2}},\penalty 10000\ \penalty 10000\ q_{1}\equiv\dfrac{\sqrt{1-p_{d}}-\sqrt{p_{d}}}{\sqrt{2}}.$ (79)
Thus, the total noisy input state is written as
$\hat{\rho}_{\text{in}}=\dfrac{1}{2^{n}}\sum\limits_{s_{1},\dots,s_{n}\in\{-,+\}^{n}}|\phi_{\text{in}}(s_{1},\dots,s_{n})\rangle\langle\phi_{\text{in}}(s_{1},\dots,s_{n})|,$ (80)
where we defined
$|\phi_{\text{in}}(s_{1},\dots,s_{n})\rangle\equiv\bigotimes_{j=1}^{n}|\phi_{s_{j}}\rangle.$ (81)
Hence, the total noisy input state is equivalent to generating a pure state $|\psi_{\text{in}}(s_{1},\dots,s_{n})\rangle$, where $(s_{1},\dots,s_{n})\in\{-,+\}^{n}$ are sampled with each probability $1/2^{n}$, i.e., uniformly.
For simplicity, as in the boson sampling case, we take the input to be $|\phi_{+}\rangle^{\otimes n}$. This assumption does not affect the simulation complexity when we consider the worst-case IQP circuit, since $|\phi_{-}\rangle=\hat{X}|\phi_{+}\rangle$ differs only by local $X$ gates, which flip $|0\rangle$ and $|1\rangle$. Adding one layer of $X$ gate into the circuit does not change the complexity, which will be rigorously shown in Sec. V.3. Therefore, the analysis for $|\phi_{+}\rangle^{\otimes n}$ as a sampled input is sufficient:
$|\phi_{\text{in}}\rangle=|\phi_{+}\rangle^{\otimes n}=\sum_{\vec{x}\in\{0,1\}^{n}}q_{0}^{n-|\vec{x}|}q_{1}^{|\vec{x}|}|\vec{x}\rangle.$ (82)
Figure 8: Pure-state sampling from mixed inputs in an IQP circuit. (a) The input $|+\rangle$ goes through the loss channel and becomes a mixed state $\hat{\tau}$. (b) We sample from an ensemble decomposition of the mixed state, drawing $|\phi_{+}\rangle$ or $|\phi_{-}\rangle$ with equal probability $1/2$.
V.2 Entanglement Rényi entropy of the output state
As discussed in Sec. II.2, an ideal IQP circuit is composed of diagonal gates, which means that a computational-basis input $|\vec{z}\rangle$ gains only a phase factor through the circuit, as $|\vec{z}\rangle\rightarrow e^{if(\vec{z})}|\vec{z}\rangle$, where $f$ is a function depending on the gates of the circuit. Using this formula, the sampled pure state input goes through the ideal circuit and transforms into the output state as follows:
$|\phi_{+}\rangle^{\otimes n}\rightarrow|\phi_{\text{out}}\rangle=\sum_{\vec{x}\in\{0,1\}^{n}}e^{if(\vec{x})}q_{0}^{n-|\vec{x}|}q_{1}^{|\vec{x}|}|\vec{x}\rangle.$ (83)
Using this form, we now obtain the upper bound of the ERE between the bipartition of this output state to determine the scaling of ERE, which can serve as a criterion for classical simulability of sampling. The system is split into two subsystems, $A$ and $B$, and subsystem $B$ is traced out to obtain the reduced density matrix of the subsystem $A$:
$\hat{\rho}_{A}$ $=\text{tr}_{B}(|\phi_{\text{out}}\rangle\langle\phi_{\text{out}}|)$ (84)
$=\sum_{\vec{x}_{A}^{(1)},\vec{x}_{A}^{(2)}\in\{0,1\}^{|A|}}\sum_{\vec{y}\in\{0,1\}^{|B|}}q_{0}^{2n-|\vec{x}_{A}^{(1)}|-|\vec{x}_{A}^{(2)}|-2|\vec{y}|}$ (85)
$\times q_{1}^{|\vec{x}_{A}^{(1)}|+|\vec{x}_{A}^{(2)}|+2|\vec{y}|}e^{i[f(\vec{x}_{A}^{(1)}:\vec{y})-f(\vec{x}_{A}^{(2)}:\vec{y})]}|\vec{x}_{A}^{(1)}\rangle\langle\vec{x}_{A}^{(2)}|,$ (86)
[…]
Here, we pick $\Phi$ as completely dephasing channel, and $\hat{\rho}_{A,\text{diag}}\equiv\Phi(\hat{\rho}_{A})$ is described as
$\hat{\rho}_{A,\text{diag}}$ $=\sum_{\vec{x}_{A}\in\{0,1\}^{|A|}}\left(\dfrac{q_{1}}{q_{0}}\right)^{2|\vec{x}_{A}|}q_{0}^{2|A|}|\vec{x}_{A}\rangle\langle\vec{x}_{A}|$ (87)
$=\bigotimes_{i\in A}\left(q_{0}^{2}|0\rangle\langle 0|_{i}+q_{1}^{2}|1\rangle\langle 1|_{i}\right).$ (88)
Hence, the Rényi entropy of $\hat{\rho}_{A,\text{diag}}$ upper-bounds the ERE of the output state of any IQP circuit. Thus, we now focus on computing the Rényi entropy of $\hat{\rho}_{A,\text{diag}}$.
Since $\hat{\rho}_{A,\text{diag}}$ is a diagonal and product form, the Rényi entropy of this state is written as
$S_{\alpha}(\hat{\rho}_{A,\text{diag}})=\dfrac{|A|}{1-\alpha}\ln(q_{0}^{2\alpha}+q_{1}^{2\alpha}).$ (89)
While it seems linearly increasing with $|A|$, if we compute the upper bound of the Rényi entropy of $\hat{\rho}_{A,\text{diag}}$, which is similarly defined, then we obtain the Rényi entropy as
$S_{\alpha}(\hat{\rho}_{B,\text{diag}})=\dfrac{|B|}{1-\alpha}\ln(q_{0}^{2\alpha}+q_{1}^{2\alpha}).$ (90)
Hence, for given subsystems $A$ and $B$, the lowest upper bound of the ERE is given as $\min(S_{\alpha}(\hat{\rho}_{A,\text{diag}}),S_{\alpha}(\hat{\rho}_{B,\text{diag}}))$. Thus, the maximum ERE is achieved when we choose $|A|=|B|$, which renders
$\max_{1\leq l\leq n-1}S_{\alpha}^{l}(|\phi_{\text{out}}\rangle)\leq\dfrac{n}{2(1-\alpha)}\ln(q_{0}^{2\alpha}+q_{1}^{2\alpha}).$ (91)
We use the inequality $\ln(q_{0}^{2\alpha}+q_{1}^{2\alpha})\leq(1-2p)^{2\alpha d}$, where $p$ denotes the noise rate per qubit per layer, to upper bound the ERE. The inequality follows from the sequence of bounds below:
$\ln(q_{0}^{2\alpha}+q_{1}^{2\alpha})\leq(1-2p_{d})^{2\alpha}=(1-2p)^{2\alpha d}.$ (92)
Here, as boson sampling, the inequalities Eq. (34), Eq. (35), and Eq. (36) are used to upper bound the term.
Recall that when the ERE of a reduced density matrix scales as $S_{\alpha}=O(\log n)$ for $0<\alpha<1$ for all bipartitions, the state is classically efficiently described by MPS. Applying this ERE scaling constraint to the upper bound of ERE, we obtain a sufficient condition for a state to be efficiently represented by MPS, which means classically simulable IQP sampling. One may see that when
$1-2p$ $\leq O\left(\left(\dfrac{\log n}{n}\right)^{1/{(2\alpha d)}}\right),$ (93)
the ERE is upper-bounded by a logarithmic scaling in $n$:
$\max_{1\leq l\leq n-1}S_{\alpha}^{l}(|\phi_{\text{out}}\rangle)$ $=\dfrac{n(1-2p)^{2\alpha d}}{2(1-\alpha)}\leq O\left(\log n\right).$ (94)
Since the depth is crucial for analyzing the classical simulability of IQP sampling, let us examine the bound of depth in terms of the noise rate $p$. If we choose $\alpha\to 1$, we can see that the sufficient condition for classical simulability is written in terms of the depth $d$ as
$d\geq O\left(\dfrac{\log n}{|\log(1-2p)|}\right).$ (95)
Hence, it shows that our MPS method is efficient up to the point where the depth scales logarithmically with the number of qubits $n$. Notably, this noise threshold is consistent with a recent result in Ref. [42], which employs a completely different method.
Figure 9: Simulation of depolarizing noise in IQP circuit. (a) Pauli $Z$ noises can commute with the very first gate of the circuit, as it is diagonal in the computational basis. (b) The noises are sampled into Pauli gates, where each $X$ and $Y$ gate occurs with probability $q$. And the input $|+\rangle$ goes through the Pauli $Z$ noise channel and becomes a mixed state $\hat{\tau}$. (c) We sample from an ensemble decomposition of the mixed state, drawing $|\phi_{+}\rangle$ or $|\phi_{-}\rangle$ with equal probability $1/2$. The $Y$ gates can be decomposed into $X$ and $Z$ gates. Here, because IQP gates are diagonal in the $Z$ basis, we absorb all $Z$-errors into the ideal IQP layer; in contrast, $X$-errors do not commute and therefore remain as the explicit noise layer.
V.3 Generalization into depolarizing noise
We now generalize this result to depolarizing noise. First, recall that depolarizing noise with parameter $p$ can be decomposed into Pauli $X$, $Y$, $Z$ noise with noise rate $q=(1-\sqrt{1-2p})/2$. Since Pauli $X$, $Y$, and $Z$ noise channels commute with each other and a Pauli $Z$ noise channel commutes with diagonal gates, as before, we can move all the Pauli $Z$ noise channels in front of all the circuits. The circuit is then given as a Pauli $Z$ noise channel followed by an IQP circuit with Pauli $X$, $Y$ noise. Thus, using the same method as before to take care of the Pauli $Z$ noise on the input state, without loss of generality, we can write the initial state, including the Pauli $Z$ noise. Let the total $Z$ noise rate over depth $d$ be $q_{d}=(1-(1-2q)^{d})/2$, then the input state
$|\phi_{\text{in}}\rangle=|\phi_{+}\rangle^{\otimes n}=\sum_{\vec{x}\in\{0,1\}^{n}}q_{0}^{n-|\vec{x}|}q_{1}^{|\vec{x}|}|\vec{x}\rangle$ (96)
has the coefficients
$q_{0}\equiv\dfrac{\sqrt{1-q_{d}}+\sqrt{q_{d}}}{\sqrt{2}},\quad q_{1}\equiv\dfrac{\sqrt{1-q_{d}}-\sqrt{q_{d}}}{\sqrt{2}}.$ (97)
We then address the remaining Pauli $X$ and $Y$ noise channels using a stochastic method. More specifically, for a Pauli $X$ ($Y$) noise channel, we probabilistically insert a Pauli $X$ ($Y$) gate with probability $q$ and the identity with probability 1q1-q. Thus, each noisy realization can be viewed as a sampled circuit containing Pauli gates, rather than a noise channel.
After sampling the Pauli gates from the Pauli noise, only $X$, Y=iZXY=-iZX, and $XY=iZ$ gates would remain in the IQP circuit. Thus, we now show how to address the remaining $X$ and $Z$ gates. First of all, $Z$ gates can be simply absorbed into the IQP circuit because they are also diagonal in the $Z$ basis. Hence, we merely have to consider a revised IQP circuit with added $Z$ gates. On the other hand, $X$ gates can also be easily addressed by the following observation. Suppose after a certain depth $t$ the quantum state is written as
$|\phi_{\text{out},t}\rangle=\sum_{\vec{x}\in\{0,1\}^{n}}e^{if_{t}(\vec{x})}q_{0}^{n-|\vec{x}|}q_{1}^{|\vec{x}|}|\vec{x}\rangle,$ (98)
where $f_{t}(\vec{x})$ is the phase due to the IQP circuit up to depth $t$. If we apply $X$ gates on this state, the quantum state transforms as
$\left[\bigotimes_{k=1}^{n}\hat{X}^{u_{k}^{(t)}}\right]|\phi_{\text{out},t}\rangle$ $=\sum_{\vec{x}\in\{0,1\}^{n}}e^{if_{t}(\vec{x})}q_{0}^{n-|\vec{x}\oplus\vec{u}^{(t)}|}q_{1}^{|\vec{x}\oplus\vec{u}^{(t)}|}|\vec{x}\rangle,$ (99)
where $u_{k}^{(t)}$ decides whether an $X$ gate is applied or not on the $k$th qubit. Thus, when we introduce the $X$ gate layers for each depth, where $u_{k}^{(t)}$ is sampled as explained above, the final state can be in the following form:
$|\phi_{\text{out}}\rangle$ $=\sum_{\vec{x}\in\{0,1\}^{n}}e^{if(\vec{x})}q_{0}^{n-|\vec{x}\oplus\vec{u}|}q_{1}^{|\vec{x}\oplus\vec{u}|}|\vec{x}\rangle,$ (100)
where $\vec{u}\equiv\sum_{t=1}^{d}\vec{u}^{(t)}$. Then, its reduced density matrix on the subsystem $A$ after removing all the off-diagonal elements is
$\hat{\rho}_{A,\text{diag}}$ $=\sum_{\vec{x}_{A}\in\{0,1\}^{|A|}}q_{0}^{2|A|-2|\vec{x}_{A}\oplus\vec{u}_{A}|}q_{1}^{2|\vec{x}_{A}\oplus\vec{u}_{A}|}|\vec{x}_{A}\rangle\langle\vec{x}_{A}|$ (101)
$=\sum_{\vec{x}_{A}\in\{0,1\}^{|A|}}q_{0}^{2|A|-2|\vec{x}_{A}|}q_{1}^{2|\vec{x}_{A}|}|\vec{x}_{A}\oplus\vec{u}_{A}\rangle\langle\vec{x}_{A}\oplus\vec{u}_{A}|.$ (102)
Since the Rényi entropy of the state in a diagonal form depends only on the coefficients, the ERE of this state is exactly the same as that we obtained from the previous section. Therefore, as in the dephasing noise case, when the circuit depth $d\geq O(\log n/|\log(1-2q)|)=O(\log n/|\log(1-2p)|)$, the IQP sampling under depolarizing noise is efficiently classically simulable. When the noise rate is small, this depth approaches $O(\frac{1}{p}\log n)$, which corresponds to the critical depth in Ref. [42].
V.4 Numerical analysis of noisy IQP sampling
We estimate the bond dimension required to simulate a dephasing-noisy IQP circuit under a fixed simulation error of $0.01$. The estimate is provided for the worst-case scenario, in which the eigenvalues of $\hat{\rho}_{A,\text{diag}}$ with $|A|=|B|=n/2$ are considered.
Figure 10: Required resource to achieve an MPS approximation error of $0.01$, as a function of qubit number and circuit depth. (a), (b) Required bond dimension and memory with a dephasing noise rate of $0.01$ per layer. Here, regardless of the depth, the unitary of the linear optics is considered to be the worst.
Since we always assume the worst-case IQP circuit regardless of the circuit depth, even with $d=1$, a highly complicated IQP circuit is considered. Due to this reason, the required bond dimension is the maximum under the same number of qubits and the same noise rate per qubit and gate. Obviously, in practice, the complexity of an IQP circuit depends on its circuit depth [54, 45, 30, 55]. Implementing a highly complicated IQP circuit in a very shallow depth is demanding in practice; thus, we expect that the bond dimension for extremely low-depth cases overestimates the complexity. In other words, for an extremely low depth, though the required bond dimension is shown to be high in our figure, the corresponding IQP sampling can be classically efficiently simulated because entanglement cannot spread to the overall qubits. Except for that case, when the circuit depth increases, the overall noise rate also increases; therefore, the complexity of the simulation decreases.
For IQP sampling, the previous work shows the simulability beyond the critical depth $d_{c}>\Theta(p^{-1}\log(p^{-1}))$ [46]. According to the reference, the phase transition for percolation occurs at depths $d_{c}\approx 33$ for noise rate $p=0.05$ and $d_{c}\approx 117$ for $p=0.02$. However, since the IQP circuit is classically simulable using MPS with a bond dimension of order $10^{4}$, which may still be possible using a large classical computer, our MPS algorithm can cover a much wider range; more specifically, ours can simulate an IQP circuit with lower depth than the critical depth of the previous work. For the calculation of simulable memory of IQP, we use the same memory resource formula as in boson sampling.
VI Discussion and conclusion
In this work, we generalized the state-of-the-art MPS method initially developed for GBS to broader classes of sampling problems, including boson sampling and IQP sampling. This generalization demonstrates that the approach is not limited to a specific platform or model but can be extended to a wide range of other sampling problems.
The proposed algorithm shows improved efficiency compared to previously proposed simulation algorithms. In addition to its efficiency, the method offers further advantages, such as tunable simulation accuracy that depends on the available computational resources. Notably, the approach remains valid across the entire range of imperfections, even in regimes where the simulation complexity increases from polynomial to exponential with respect to system parameters. Hence, we expect that our new algorithm enables us to simulate a lossy boson sampling and noisy IQP sampling in a practical regime and can be used as a benchmark for quantum advantage experiments.
Furthermore, we provide explicit resource estimates by analyzing the required bond dimension to achieve a target simulation noise rate. This quantitative analysis enables informed decisions about simulation feasibility. In particular, for circuits corresponding to an experimentally implemented unitary, we conduct a detailed simulability analysis that aligns closely with a realistic experimental setup. These results enable a direct comparison between theoretical predictions and experimental outcomes, thereby enhancing the practical relevance of our approach and the usage of our algorithm as a benchmark for future experimental implementations.
We remark that although our method for lossy boson sampling covers Fock-state and cat-state inputs, whether it can be further generalized to an arbitrary input state is still open. Also, while we assumed a conventional IQP circuit [19], a recent study in Ref. [42] extended the IQP circuit to be augmented by CNOT gates. Our method does not immediately extend to those IQP circuits yet. It is open to improve our algorithm to cover a broader class of quantum circuits. Also, we assumed a lossy channel or Pauli noise for boson sampling and IQP sampling, respectively. Generalizing our method to cover more general noise channels would be an important open question.
Acknowledgements.
S.P. and C.O. were supported by the National Research Foundation of Korea Grants (No. RS-2024-00431768 and No. RS-2025-00515456) funded by the Korean government (Ministry of Science and ICT (MSIT)) and the Institute of Information & Communications Technology Planning & Evaluation (IITP) Grants funded by the Korea government (MSIT) (No. IITP-2025-RS-2025-02283189 and IITP-2025-RS-2025-02263264).
Appendix A Analysis of the required bond dimension for classical simulation of boson sampling
This section describes the algorithm for estimating the required bond dimension in the classical simulation of boson sampling, which is used for Fig. 6. Recall that the eigenvalues of the reduced density matrix for each $j$th tensor product of Eq. (30) are given by
$c_{j0}\equiv\dfrac{1}{4}\left(2+\sqrt{2\eta^{2}\cos(4\theta_{j})-2\eta^{2}+4}\right),\qquad c_{j1}\equiv\dfrac{1}{4}\left(2-\sqrt{2\eta^{2}\cos(4\theta_{j})-2\eta^{2}+4}\right),$ (103)
where $1\geq c_{j0}\geq 1/2\geq c_{j1}\geq 0$. Hence, the eigenvalues of the total reduced density matrix are expressed as:
$p(\vec{k})\equiv\prod_{\vec{k}\in\{0,1\}^{N}}c_{jk_{j}}=\prod_{j=1}^{N}c_{j0}^{1-k_{j}}c_{j1}^{k_{j}}.$ (104)
Hence,
$\log p(\vec{k})$ $=\sum\limits_{j=1}^{N}\left[(1-k_{j})\log c_{j0}+k_{j}\log c_{j1}\right]=\sum\limits_{j=1}^{N}\log c_{j0}+\sum\limits_{j=1}^{N}k_{j}\log(\dfrac{c_{j1}}{c_{j0}})=C+\sum\limits_{j=1}^{N}k_{j}\log r_{j},$ (105)
where we define $C\equiv\sum_{j=1}^{N}\log c_{j0}$ and $r_{j}\equiv c_{j1}/c_{j0}<1$. Note that as $r_{j}$ decreases, $p(\vec{k})$ becomes a smaller eigenvalue and thus less significant.
We construct the list of $\log p(\vec{k})$ iteratively. Initialize $L=\{0\}$, and for each $j$ update
$L\leftarrow L\,\cup\,\{\;\ell+\log r_{j}\;:\;\ell\in L\;\}.$ (106)
Since $\log r_{j}<0$, we can prune partial sums during the construction: for any $t<N$, if a partial value $\sum_{j=1}^{t}k_{j}^{*}\log r_{j}$ is below the threshold, then every $\vec{k}$ whose first $t$ entries coincides with $\vec{k}^{*}$ will also be below the threshold; hence we discard that entire branch from $L$. This pruning removes eigenvalues below the threshold.
By choosing the threshold so that the cumulative weight of the retained eigenvalues is sufficiently close to the target accuracy, say $99\%$, and then counting the eigenvalues that exceed the threshold, we obtain an estimate of the required bond dimension.
Appendix B Commutation relation of splitted creation and annihilation operators
In this Appendix, we provide a detailed proof for the statement that the commutation relation of $\hat{B}_{\text{u},j}$ approximately holds for Haar-random unitary matrices $U$. Recall that the commutation relation is written as
$\left[\hat{B}_{\text{u},j},\hat{B}_{\text{u},k}^{\dagger}\right]$ $=\left[\dfrac{\sum\limits_{m=1}^{l}U_{jm}\hat{a}_{m}}{\cos\theta_{j}},\dfrac{\sum\limits_{n=1}^{l}U_{kn}^{*}\hat{a}_{n}^{\dagger}}{\cos\theta_{k}}\right]=\dfrac{\sum\limits_{m=1}^{l}U_{jm}U_{km}^{*}}{\sqrt{\sum\limits_{m=1}^{l}|U_{jm}|^{2}}\sqrt{\sum\limits_{n=1}^{l}|U_{kn}|^{2}}},$ (107)
for $1\leq j,k\leq N$. We will focus on the case that $l=M/2$ because this case renders the maximum entanglement entropy.
First of all, the commutation relation for $j=k$ holds trivially. Thus, we focus on the case that $j\neq k$. We now show that the commutation relation with $j\neq k$ is small with high probability over the Haar-random unitary $U$. To do that, we first compute the first and second moments of the related terms. Let us first invoke some identities from the Weingarten calculus:
$\mathbb{E}[U_{im}U_{jm}^{*}]=\frac{\delta_{ij}}{M},\quad\mathbb{E}[|U_{im}U_{jn}^{*}|^{2}]=\frac{1+\delta_{ij}}{M^{2}-1}-\frac{\delta_{jk}+\delta_{mn}}{M(M^{2}-1)},$ (108)
Let us first focus on the numerator. The first and second moments of the numerator are given by
$\mathbb{E}\left[\sum_{m=1}^{l}U_{jm}U_{km}^{*}\right]=0,\quad\mathbb{E}\left[\left|\sum_{m=1}^{l}U_{jm}U_{km}^{*}\right|^{2}\right]=\dfrac{l(l+1)}{M^{2}(M+1)}-\dfrac{l^{2}}{M^{2}}=\dfrac{l(M-l)}{M^{2}(M+1)}.$ (109)
Hence, the variance of the numerator is the same as the second moment. Therefore, using the Chebyshev inequality,
$\Pr[\left|\sum_{m=1}^{l}U_{jm}U_{km}^{*}\right|\leq\epsilon]\leq 1-\frac{1}{\epsilon^{2}}\dfrac{l(M-l)}{M^{2}(M+1)}.$ (110)
We similarly investigate the denominator
$\mathbb{E}\left[\sum_{m=1}^{l}|U_{jm}|^{2}\right]=\frac{l}{M},\quad\mathbb{E}\left[\left|\sum_{m=1}^{l}|U_{jm}|^{2}\right|^{2}\right]=\frac{l(l+1)}{M(M+1)},$ (111)
and thus the variance is $l(M-l)/[M^{2}(M+1)]$. Hence, by the Chebyshev inequality,
$\Pr[\left|\sum_{m=1}^{l}|U_{jm}|^{2}-\frac{l}{M}\right|\leq\epsilon]\leq 1-\frac{1}{\epsilon^{2}}\dfrac{l(M-l)}{M^{2}(M+1)}.$ (112)
Hence, using the union bound, we prove that
$\Pr[\left|\sum_{m=1}^{l}|U_{jm}|^{2}-\frac{l}{M}\right|\leq\epsilon\penalty 10000\ \text{and}\penalty 10000\ \left|\sum_{m=1}^{l}|U_{km}|^{2}-\frac{l}{M}\right|\leq\epsilon\penalty 10000\ \text{and}\penalty 10000\ \left|\sum_{m=1}^{l}U_{jm}U_{km}^{*}\right|\leq\epsilon]\leq 1-\frac{3}{\epsilon^{2}}\dfrac{l(M-l)}{M^{2}(M+1)}.$ (113)
Since a similar bound must hold for all $1\leq j,k\leq N$, by using the union bound once more,
$\Pr[\left|\sum_{m=1}^{l}|U_{jm}|^{2}-\frac{l}{M}\right|\leq\epsilon\penalty 10000\ \text{and}\penalty 10000\ \left|\sum_{m=1}^{l}|U_{km}|^{2}-\frac{l}{M}\right|\leq\epsilon\penalty 10000\ \text{and}\penalty 10000\ \left|\sum_{m=1}^{l}U_{jm}U_{km}^{*}\right|\leq\epsilon\penalty 10000\ \text{for all}\penalty 10000\ 1\leq j,k\leq N]$ (114)
$\leq 1-\frac{3N(N-1)}{8\epsilon^{2}(M+1)},$ (115)
where we set $l=M/2$ as assumed. Hence, by choosing $M=\omega(N^{2})$ for any $\epsilon$,
$\left|\left[\hat{B}_{\text{u},j},\hat{B}_{\text{u},k}^{\dagger}\right]\right|=\dfrac{\sum\limits_{m=1}^{l}U_{jm}U_{km}^{*}}{\sqrt{\sum\limits_{m=1}^{l}|U_{jm}|^{2}}\sqrt{\sum\limits_{n=1}^{l}|U_{kn}|^{2}}}\leq\frac{\epsilon}{M/2-\epsilon}=O(1/M),$ (116)
which proves our assumption.